OpenStudy (anonymous):

Find the equation of the parabola if vertex at (2, 5), axis parallel to the x-axis, passing through (0, 4)

OpenStudy (anonymous):

@jim_thompson5910 help please :)

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

not far at all... do we use the equation y = (x - h)2 + k

jimthompson5910 (jim_thompson5910):

that's if the line of symmetry is parallel to the y axis

OpenStudy (anonymous):

or there is anothe one we learnt at school which is (y-k)^2 = 4a(x-h)

jimthompson5910 (jim_thompson5910):

that's the one we'll use

jimthompson5910 (jim_thompson5910):

the vertex is (2,5), so h = 2 k = 5 plug those values in

OpenStudy (anonymous):

(y-5)^2 = 4a(x-2)

jimthompson5910 (jim_thompson5910):

you also know that (0,4) is on the parabola so plug x = 0 and y = 4 in as well

jimthompson5910 (jim_thompson5910):

then solve for 'a'

OpenStudy (anonymous):

a = 1/8

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

negative? a = -1/8

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

since a = -1/8, this means 4a = 4(-1/8) = -4/8 = -1/2

jimthompson5910 (jim_thompson5910):

so (y-5)^2 = 4a(x-2) becomes (y-5)^2 = -1/2(x-2)

OpenStudy (anonymous):

Thanks again @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

you're welcome